# Frame In-N-Out: Unbounded Controllable Image-to-Video Generation

### v1

@@ Abstract
Controllability, temporal coherence, and detail synthesis remain the most critical challenges in video generation. In this paper, we focus on a commonly used yet underexplored cinematic technique known as Frame In and Frame Out. Specifically, starting from image-to-video generation, users can control the objects in the image to naturally leave the scene or provide breaking new identity references to enter the scene, guided by user-specified motion trajectory. To support this task, we introduce a new dataset curated semi-automatically, a comprehensive evaluation protocol targeting this setting, and an efficient identity-preserving motion-controllable video Diffusion Transformer architecture. Our evaluation shows that our proposed approach significantly outperforms existing baselines.
@@ 1 Introduction
We formalize Frame In and Frame Out as an image-to-video generation setting. Starting from the first frame image, our goal is to create a model that, given an explicit motion trajectory, can (i) control an existing object in the frame completely outside the visible bounds of the first frame and subsequently bring it back while preserving fidelity and integrity, which is defined as Frame Out, and (ii) allow a new identity (ID) object (e.g., humans, vehicles, animals) to enter the scene plausibly—whether from the sides or from above, which is defined as Frame In. Contemporary motion-controlled image-to-video generation architecture design like [82, 15, 62, 87, 78, 45, 53] needs spatiotemporal pixel-aligned trajectory signal to the first frame, treating the image border as an immovable “wall”, which we can see from Fig. 4. To transcend this border-bounded constraint, we extend the conditioning control beyond the first frame region size to an unbounded canvas region. Conditioning signals are applied over this enlarged canvas, enabling the object to move out of the first frame region or using the canvas region to prepare the entrance of a new identity (ID) reference while maintaining temporal and spatial coherence.
The overall design of the Frame In-N-Out framework is under study. A key challenge lies in the absence of existing training datasets that explicitly capture Frame In and Frame Out dynamics. To address this, we redesign the data curation pipeline from scratch. This includes identifying track-worthy objects, improving tracking reliability, and defining suitable bounding boxes that partition the first frame from the extended canvas region. A new video Diffusion Transformer [77, 41] is needed to integrate multiple conditions that are either spatiotemporal pixel-aligned [33, 78, 68] (e.g., motion), unaligned [34, 13] (e.g., identity), and more importantly, unbounded canvas. As Frame In-N-Out represents a new task in the video generation domain, we meticulously curated a dedicated evaluation system. This includes constructing benchmark datasets for both Frame In and Frame Out scenarios and revising traditional tracking, and identity-preserving metrics to reflect the unique demands of this problem setting. We believe this formulation can inspire future research into more expressive and application-driven conditioning research in video generation.
@@
  • We define and curate a training dataset for the Frame In and Frame Out pattern recognition. The processed metadata will be released.
  • We propose an efficient controllable video Diffusion Transformer that unifies spatiotemporal pixel-aligned motion, unaligned identity reference, and our proposed unbounded canvas conditions at the minimum cost but with stable generative results.
  • We provide an evaluation system with a testing dataset and adjusted traditional metrics for the Frame In and Frame Out scenarios. We believe our insight and area focus has broad prospects in scenarios such as the film industry or advertising production.
@@ 2 Related Works
CogVideoX [77]
MotionCtrl [68]
DragAnything [70]
Image Conductor [33]
ToRA [87]
ConsisID [80]
SkyReals-A2 [13]
Phantom [36]
@@
Base Condition in Video Generation. While numerous conditioning strategies have been proposed for video generation, some base conditioning exerts a fundamental influence on the eventual generation quality and choice of the base model. Broadly, these can be categorized into three primary paradigms: Text-to-Video [5, 18, 9, 57], Image-to-Video [7, 77, 57], and Video-to-Video [32, 76, 38, 54, 3]. Vanilla text-to-video tasks involve generation from sparse conditioning signals, where no pixel-level guidance is needed. The model must rely entirely on language prompts to imagine and synthesize all visual content, including scene layout, motion, and object appearance. In contrast, Image-to-Video assumes the presence of a single reference frame as the first frame, from which the entire video must follow. This requires all subsequent frames to remain aligned with the initial spatial content even if additional conditions like text prompts are provided.
Controllable Video Generation. Controllable video generation refers to the task of extending pre-trained video generation models by modifying their architecture to incorporate one or more additional conditions beyond the original text, image, or video inputs. Since image diffusion models, researchers have explored a wide range of conditioning signals. This including sketches [72], human pose [28], low-quality images [69] for restoration, masked images for inpainting [25, 91], outpainting [85, 10], and editing [79, 24]. In the video generation domain, temporal-oriented challenges include interpolation between the first and last frame [23, 58], motion control [68, 73, 75, 78, 89, 67, 62], camera control [71, 68, 17, 67, 3, 73], and long-range history-guided generation [46, 81]. Further, as a flexible condition, identity (ID) reference is also broadly studied in both image and video side, like PhotoMaker [34], InstantID [66], ConsisID [80], and Concat-ID [90]. Building upon this foundation, there has been a growing trend toward elements-to-video generation, where not only identity reference images but also the first frame can serve as an individual compositional element, like Phantom [36], and SkyReels-A2 [13]. The conditioning comparisons can be found in Tab. 1.
@@ 3 Problem Definition
This paper focuses on solving unbounded controllable image-to-video generation. Specifically, we concrete the problem into a specific task in the cinematic domain: Frame In and Frame Out. Our controllable video generation targets at the intersection of four control signals: (1) first frame image $\textit{I}_{0}$ and text prompt $\mathbf{y}$ as the foundation condition, (2) a canvas area expansion bounding box setting $\mathbf{B}_{\text{canvas}}$ that is composed of top-left and bottom-right pixels enlarge setting, (3) motion trajectory $\mathbf{c}_{\text{trajs}}$ for an existing object in the first frame or a new identity (ID) to introduce, and (4) an optional identity reference image $\mathbf{f}$ (e.g., a human, vehicle, animal, balloon, etc.). In the Frame Out case, we don’t need an ID reference provided, but it is mandatory in the Frame In case. Eventually, our video Diffusion Model generates $N$ number of latent frames that strictly follow all conditions, aiming to learn a conditional joint distribution $p_{\theta}(\textit{I}_{1},...,\textit{I}_{N}|\textit{I}_{0},\mathbf{y},\mathbf{B}_{\text{canvas}},\mathbf{c}_{\text{trajs}},\mathbf{f})$.
@@ 4 Dataset Curation and In-N-Out Pattern Recognition
Basic Curation. Our basic curation consists of the following steps. (1) Metadata filtering: we first selected videos based on the metadata attributes such as duration, resolution, and aspect ratio. (2) Image-level filtering: for each video, we randomly sample two frames and filter out low-quality videos using automated image quality assessment [49] and aesthetic assessment [50]. We additionally apply image complexity assessment [14] to exclude both overly simplistic and excessively complex videos, which are known to hinder learning [60]. Videos with excessive overlaid text are also filtered using an OCR detector [2]. (3) Video-level filtering: we remove multi-scene videos using scene cut detection from TransNet V2 [47], and discard videos with significant camera motion (e.g., rotation, translation, or focal changes) based on motion estimation from CUT3R [65], focusing on object-centric motion with least impact from camera movement. (4) Automatic Captioning: to obtain high-quality paired text prompts, we discard dataset-provided captions and generate new ones by QWen2.5-32B-VL-Instruct [74].
@@
Identity of Interest. Random point-based tracking, like optical flow [68, 87], does not provide semantic meaning to each point; i.e. each point cannot strictly correspond to an identity. Thus, before applying the tracking model, we apply panoptic segmentation [30] with OneFormer [22] to classify and then segment all objects in the image. We observe that videos with ideal Frame In and Frame Out patterns usually come with multiple relevant start frames across a video. Thus, we select 3 starter frames at the duration 0%, 35%, and 70% of the full video length. This strategy alleviates the dataset scarcity in the later stage. To get the clearest and the least compressed frames available, we apply I-Frame extraction from traditional video compression [44, 60] and choose the closest I-Frame as the official starter frame. These 3 starter frames will be taken as the first frame for image-to-video generation and also execute the panoptic segmentation. Panoptic Segmentation from OneFormer will classify 133 classes based on the COCO dataset [35]. We manually define 22 classes of them as motionable objects that could be objects of interest in the following tracking annotation. Our purpose is to filter out static objects, like trees, houses, and sightseeing, which are not ideal as a tracking target. Meanwhile, based on the size of segmentation masks, we filter both small and over-sized identity objects. Further, inspired by [62], we apply K-means to get an even distribution of points from the segmentation mask.
@@
Cycle Tracking: With objects of interest and even distributed points, we apply tracking from CoTracker3 [27]. However, tracking can be unstable and inaccurate, especially on fast-moving objects. To the best of our capability to provide the most accurate tracking trajectories without human correction, we take advantage of back-tracking functionality from CoTracker3. After the regular forward track, we back-track from the end position of points in the last frame to the first frame. If the error between the initial position and the back-tracked position is larger than a preset threshold, we filter out those points. Less but accurate points are more helpful for motion-controllable video generation training. We perform dense tracking for 24 FPS to ensure the highest accuracy though we might only need 12 FPS of information in the training. In the end, we sort and filter small and extremely high-motion cases to avoid static objects and over-fast movements.
@@
Frame Out cases are instances where the object is initially partially or completely located inside the bounding box and subsequently moves entirely outside the bounding box in at least one frame of the video. Re-entry into the box is allowed for robust training and diverse user cases. In contrast, Frame In cases require the object to be completely outside the bounding box in the first frame, with no pixel overlap. This ensures that the ID reference has no overlap with the first frame in the training and we could condition breaking new information to the image-to-video generation. Next, in the following frames, a sufficient fraction of the object must come to the in-box region to be considered as a qualified Frame In scenario. For valid Frame In cases, we employ SAM2 [43] to extract the object mask and store the corresponding cropped ID reference image. The SAM2 mask is also used to further filter out inaccurate tracking points.
@@ 5 Frame In-N-Out Architecture
Figure 3: Main Architecture. Our video Diffusion Transformer embraces the first frame with canvas expansion, motion, identity reference, and text prompt as the conditions for video generation.
@@ 5.1 Base Architecture with Flexible Resolution Training
The base video Diffusion Transformer architecture we consider is CogVideoX-I2V [77], a relatively small-scale model (5 billion parameters) compared to tens of billions of parameter models [58, 31, 37, 16]. We believe that our contribution is scalable to most video Diffusion Transformer considering the similarity in the architecture [41]. The first frame $\mathit{I}_{0}$ is conditioned by doing channel-wise concatenation with the noisy latent $\mathbf{Z}$ before entering the transformer blocks:
@@
Training in default resolution of 480$\times$720 is computation-heavy. Hence, we take advantage of the nature of the absolute position embedding and rotary position embedding [48] (RoPE) by training in lower resolutions 384$\times$480 and then testing on versatile resolutions that the user uses. To do so, inspired by  [^1] https://github.com/aigc-apps/VideoX-Fun, we first modify the absolute position embedding to support versatile resolution inputs by applying a trilinear interpolation on the learned fixed absolute embeddings to the target latent size. For RoPE, we inject the current resolution grid size to create a new position embedding each time.
@@ 5.2 Motion Control
We first convert all spatiotemporal trajectory coordinates into a pixel marking in image forms $\mathbf{c}_{trajs}$. Since we use panoptic segmentation in the dataset curation, we have rich and accurate semantic meaning for each tracking point. Hence, we provide different objects with different color markings to promote the model learning semantic relationship. For the same object with multiple tracking points, they share the same color. Our trajectory point represents the spatiotemporal pixel-aligned state of the object which is different from optical flow-based motion vector representation as in previous motion control works [68, 59, 87].
The motion conditioning has various solutions in the literature. ControlNet-like [82, 15, 62], cross-attention based [87, 78], or extra training tokens [26] methods are all computational expensive for motion conditioning. For pixel-aligned conditions, we prefer to apply a direct, efficient, and natural solution, which is by channel-wise concatenating VAE-encoded motion images. To be specific, we encode motion images like regular RGB images by the pre-trained 3D VAE encoder $\mathcal{E}$; thus, the latent size of the motion is perfectly aligned with the first frame latent and noisy latent $\mathbf{Z}$. Then, we do channel-wise concatenation to the motion latent as the following:
@@
However, channel-wise concatenation increases the input channel number for the first projector of the Diffusion Transformer. Hereby, inspired by Marigold [29], we first zero-initialized the projector with the new input channel number setting and then filled in the overlap weights with the pre-trained weights available to decrease the training gap.

### v2

@@ Abstract
Controllability, temporal coherence, and detail synthesis remain the most critical challenges in video generation. In this paper, we focus on a commonly used yet underexplored cinematic technique known as Frame In and Frame Out. Specifically, starting from image-to-video generation, users can control the objects in the image to naturally leave the scene or provide breaking new identity references to enter the scene, guided by a user-specified motion trajectory. To support this task, we introduce a new dataset that is curated semi-automatically, an efficient identity-preserving motion-controllable video Diffusion Transformer architecture, and a comprehensive evaluation protocol targeting this task. Our evaluation shows that our proposed approach significantly outperforms existing baselines.
@@ 1 Introduction
We formalize Frame In and Frame Out as an image-to-video generation setting. Starting from the first frame image (as shown in Fig. 1), our goal is to create a model that, given an explicit motion trajectory, can (i) control an existing object in the frame completely outside the visible bounds of the first frame and subsequently bring it back while preserving fidelity and integrity, which is defined as Frame Out, and (ii) allow a new identity (ID) object (e.g., humans, vehicles, animals) to enter the scene plausibly—whether from the sides or from above, which is defined as Frame In. Contemporary motion-controlled image-to-video generation architecture design like zhang2023adding ; gu2025diffusion ; wang2024levitor ; zhang2024tora ; yin2023dragnuwa ; shi2024motionstone ; tanveer2024motionbridge needs spatiotemporal pixel-aligned trajectory signal to the first frame, treating the image border as an immovable “wall”, which we can see from Fig. 4. To transcend this border-bounded constraint, we extend the conditioning control beyond the first frame region size to an unbounded canvas region. Conditioning signals are applied over this enlarged canvas, enabling the object to move out of the first frame region or using the canvas region to prepare the entrance of a new identity (ID) reference while maintaining temporal and spatial coherence.
The overall design of the Frame In-N-Out framework is under study. A key challenge lies in the absence of existing training datasets that explicitly capture Frame In and Frame Out dynamics. To address this, we redesign the data curation pipeline from scratch. This includes identifying track-worthy objects, improving tracking reliability, and defining suitable bounding boxes that partition the first frame from the extended canvas region. A new video Diffusion Transformer yang2024cogvideox ; peebles2023scalable is needed to integrate multiple conditions that are either spatiotemporal pixel-aligned li2025image ; yin2023dragnuwa ; wang2024motionctrl (e.g., motion), unaligned li2024photomaker ; fei2025skyreels (e.g., identity), and more importantly, unbounded canvas. As Frame In-N-Out represents a new task in the video generation domain, we meticulously curated a dedicated evaluation system. This includes constructing benchmark datasets for both Frame In and Frame Out scenarios and revising traditional tracking, and identity-preserving metrics to reflect the unique demands of this problem setting. We believe this formulation can inspire future research into more expressive and application-driven conditioning research in video generation.
@@
  • We define and curate a training dataset for the Frame In and Frame Out pattern recognition. The pre-processing code and metadata will be released.
  • We propose an efficient controllable video Diffusion Transformer that unifies spatiotemporal pixel-aligned motion, pixel-unaligned identity reference, and our proposed unbounded canvas conditions in one model.
  • We provide an evaluation system for the Frame In and Frame Out scenarios and showcase that ours outperform other baselines. Our insight and area focus have broad prospects in scenarios such as the film industry or advertising production.
@@ 2 Related Works
CogVideoX yang2024cogvideox
MotionCtrl wang2024motionctrl
DragAnything wu2024draganything
Image Conductor li2025image
ToRA zhang2024tora
ConsisID yuan2024identity
SkyReels-A2 fei2025skyreels
Phantom liu2025phantom
@@
Base Condition in Video Generation. While numerous conditioning strategies have been proposed for video generation, some base conditioning exerts a fundamental influence on the eventual generation quality and choice of the base model. Broadly, these can be categorized into three primary paradigms: Text-to-Video bao2024vidu ; henschel2024streamingt2v ; chen2024videocrafter2 ; wan2025wan , Image-to-Video blattmann2023stable ; yang2024cogvideox ; wan2025wan , and Video-to-Video ku2024anyv2v ; yang2023rerender ; mou2024revideo ; tu2025videoanydoor ; bai2025recammaster ; zhang2025flexiact . Vanilla text-to-video task involves generation from sparse conditioning signals, where no pixel-level guidance is needed. The model must rely entirely on language prompts to imagine and synthesize all visual content, including scene layout, motion, and object appearance. In contrast, Image-to-Video assumes the presence of a single reference frame as the first frame, from which the entire video must follow. This requires all subsequent frames to remain aligned with the initial spatial content, even if additional conditions like text prompts are provided.
Controllable Video Generation. Controllable video generation refers to the task of extending pre-trained video generation models by modifying their architecture to incorporate one or more additional conditions beyond the original text, image, or video inputs. Since the appearance of the image diffusion models, researchers have explored a wide range of conditioning signals. This including sketches xing2024tooncrafter , human pose karras2023dreampose , low-quality images wu2024one for restoration, masked images for inpainting ju2024brushnet ; zhou2023propainter , outpainting zhang2024continuous ; chen2024follow , and editing yu2025objectmover ; jiang2025vace . In the video generation domain, temporal-oriented challenges include interpolation between the first and last frame jain2024video ; wang2025wan , motion control wang2024motionctrl ; xing2025motioncanvas ; yang2024direct ; yin2023dragnuwa ; zheng2025vidcraft3 ; wang2024objctrl ; wang2024levitor , camera control xiao20243dtrajmaster ; wang2024motionctrl ; he2024cameractrl ; wang2024objctrl ; bai2025recammaster ; xing2025motioncanvas , and long-range history-guided generation song2025history ; zhang2025packing . Further, as a flexible condition, identity (ID) reference is also broadly studied in both image and video side, like PhotoMaker li2024photomaker , InstantID wang2024instantid , ConsisID yuan2024identity , and Concat-ID zhong2025concat . Building upon this foundation, there has been a growing trend toward elements-to-video generation, where not only identity reference images but also the first frame can serve as an individual compositional element, like Phantom liu2025phantom , and SkyReels-A2 fei2025skyreels . The conditioning comparisons can be found in Tab. 1.
@@ 3 Problem Definition
This paper focuses on solving unbounded controllable image-to-video generation. Specifically, we concretize the problem into a specific task in the cinematic domain: Frame In and Frame Out. Our controllable video generation targets at the intersection of four control signals: (1) first frame image $\textit{I}_{0}$ and text prompt $\mathbf{y}$ as the foundation condition, (2) a canvas area expansion bounding box setting $\mathbf{B}_{\text{canvas}}$ that is composed of top-left and bottom-right pixels expansion amount, (3) motion trajectory $\mathbf{c}_{\text{trajs}}$ for an existing object in the first frame or a new identity (ID) to introduce, and (4) an optional identity reference image $\mathbf{f}$ (e.g., a human, vehicle, animal, balloon, etc.). In the Frame Out case, we don’t need an ID reference provided, but it is mandatory in the Frame In case. Eventually, our video Diffusion Model generates $N$ number of latent frames that strictly follow all conditions, aiming to learn a conditional joint distribution $p_{\theta}(\textit{I}_{1},...,\textit{I}_{N}|\textit{I}_{0},\mathbf{y},\mathbf{B}_{\text{canvas}},\mathbf{c}_{\text{trajs}},\mathbf{f})$.
@@ 4 Dataset Curation and In-N-Out Pattern Recognition
Basic Curation. Our basic curation consists of the following steps. (1) Metadata filtering: we first selected videos based on the metadata attributes such as duration, resolution, and aspect ratio. (2) Image-level filtering: for each video, we randomly sample two frames and filter out low-quality videos using automated image quality assessment su2020blindly and aesthetic assessment talebi2018nima . We additionally apply image complexity assessment feng2022ic9600 to exclude both overly simplistic and excessively complex videos, which are known to hinder learning wang2024apisr . Videos with excessive overlaid text are also filtered using an OCR detector baek2019character . (3) Video-level filtering: we remove multi-scene videos using scene cut detection from TransNet V2 soucek2024transnet , and discard videos with significant camera motion (e.g., rotation, translation, or focal changes) based on motion estimation from CUT3R wang2025continuous , focusing on object-centric motion with least impact from camera movement. (4) Automatic Captioning: to obtain high-quality paired text prompts, we discard dataset-provided captions and generate new ones by QWen2.5-32B-VL-Instruct yang2024qwen2 .
@@
Identity of Interest. Random point-based tracking, like optical flow wang2024motionctrl ; zhang2024tora , does not provide semantic meaning to each point; i.e. each point cannot strictly correspond to an identity. Thus, before applying the tracking model, we apply panoptic segmentation kirillov2019panoptic with OneFormer jain2023oneformer to classify and then segment all objects in the image. We observe that videos with ideal Frame In and Frame Out patterns usually come with multiple relevant start frames across a video. Thus, we select 3 starter frames at the duration 0%, 35%, and 70% of the full video length. This strategy alleviates the dataset scarcity in the later stage. To get the clearest and the least compressed frames available, we apply I-Frame extraction from traditional video compression schwarz2007overview ; wang2024apisr and choose the closest I-Frame as the official starter frame. These 3 starter frames will be taken as the first frame for image-to-video generation and also execute the panoptic segmentation. Panoptic Segmentation from OneFormer will classify 133 classes based on the COCO dataset lin2014microsoft . We manually define 22 classes of them as motionable objects that could be objects of interest in the following tracking annotation. Our purpose is to filter out static objects, like trees, houses, and sightseeing, which are not ideal as a tracking target. Meanwhile, based on the size of segmentation masks, we filter both small and over-sized identity objects. Further, inspired by wang2024levitor , we apply K-means to get an even distribution of points from the segmentation mask.
@@
Cycle Tracking: With objects of interest and even distributed points, we apply tracking from CoTracker3 karaev2024cotracker3 . However, tracking can be unstable and inaccurate, especially on fast-moving objects. To the best of our capability to provide the most accurate tracking trajectories without human correction, we take advantage of back-tracking functionality from CoTracker3. After the regular forward track, we back-track from the end position of points in the last frame to the first frame. If the error between the initial position and the back-tracked position is larger than a preset threshold, we filter out those points. Fewer but accurate points are more helpful for motion-controllable video generation training. In the end, we sort and filter small and extremely high-motion cases to avoid static objects and over-fast movements.
@@
Frame Out cases are instances where the object is initially partially or completely located inside the bounding box and subsequently moves entirely outside the bounding box in at least one frame of the video. Re-entry into the box is allowed for robust training and diverse user cases. In contrast, Frame In cases require the object to be completely outside the bounding box in the first frame, with no pixel overlap. This ensures that the ID reference has no overlap with the first frame in the training and we could condition breaking new information to the image-to-video generation. Next, in the following frames, a sufficient fraction of the object must come to the in-box region to be considered as a qualified Frame In scenario. For valid Frame In cases, we employ SAM2 ravi2024sam to extract the object mask and store the corresponding cropped ID reference image. The SAM2 mask is also used to further filter out inaccurate tracking points.
@@ 5 Frame In-N-Out Architecture
Figure 3: Main Architecture. Our video Diffusion Transformer embraces the first frame with canvas expansion, motion trajectories, identity reference, and text prompt as conditions for video generation.
@@ 5.1 Base Architecture with Flexible Resolution Training
The base video Diffusion Transformer architecture we consider is CogVideoX-I2V yang2024cogvideox , a relatively small-scale model (5 billion parameters) compared to tens of billions of parameter models wang2025wan ; kong2024hunyuanvideo ; ma2025step ; hacohen2024ltx . We believe that our contribution is scalable to most video Diffusion Transformer considering the similarity in the architecture peebles2023scalable . The first frame $\mathit{I}_{0}$ is conditioned by doing channel-wise concatenation with the noisy latent $\mathbf{Z}$ before entering the transformer blocks:
@@
[…]
@@ 5.2 Motion Control
We first convert all spatiotemporal trajectory coordinates into a pixel marking in image forms $\mathbf{c}_{trajs}$. Since we use panoptic segmentation in the dataset curation, we have rich and accurate semantic meaning for each tracking point. Hence, we provide different objects with different color markings to promote the model learning semantic relationship. For the same object with multiple tracking points, they share the same color. Our trajectory point represents the spatiotemporal pixel-aligned state of the object which is different from optical flow-based motion vector representation as in previous motion control works wang2024motionctrl ; wang2024language ; zhang2024tora .
The motion conditioning has various solutions in the literature. ControlNet-like zhang2023adding ; gu2025diffusion ; wang2024levitor , cross-attention based zhang2024tora ; yin2023dragnuwa , or extra training tokens ju2025fulldit methods are all computational expensive for motion conditioning. For pixel-aligned conditions, we prefer to apply a direct, efficient, and natural solution, which is by channel-wise concatenating VAE-encoded motion images. To be specific, we encode motion images like regular RGB images by the pre-trained 3D VAE encoder $\mathcal{E}$; thus, the latent size of the motion is perfectly aligned with the first frame latent and noisy latent $\mathbf{Z}$. Then, we do channel-wise concatenation to the motion latent as the following:
@@
However, channel-wise concatenation increases the input channel number for the first projector of the Diffusion Transformer. Hereby, inspired by Marigold ke2024repurposing , we first zero-initialized the projector with the new input channel number setting and then filled in the overlap weights with the pre-trained weights available to decrease the training gap.